OpenStudy (anonymous):

power reduction identities express in terms of cosines to the power of 1 2sin^6x

OpenStudy (anonymous):

\[2\sin ^{6}x\]

OpenStudy (anonymous):

\[2\left( \sin ^{2}x \right) \ \left( \sin ^{2}x \right)\left( \sin ^{2}x \right)\]